OpenStudy (anonymous):

Find the remainder if P(x)=x^3+17 is divided by (x+2).

OpenStudy (anonymous):

9 ?

OpenStudy (anonymous):

The answer is 9, but I dont understand how to get it.

OpenStudy (anonymous):

@DarthTony

Directrix (directrix):

@blahblah_who_cares Have you studied synthetic division or the Remainder Theorem?

OpenStudy (anonymous):

I have.

Directrix (directrix):

Remainder Theorem?

OpenStudy (anonymous):

No. Just synthetic. And I might have learned Remainder but I dont recall anything like it. .

OpenStudy (dumbcow):

for (x+2) evaluate f(-2) to get remainder

Directrix (directrix):

We can do either here. I will post the Remainder Theorem so that you can tell me if you've ever seen it. Regardless, we will do the problem but will use the method you have studied. Tell me which method after you look at this remainder theorem. Remainder Theorem If a polynomial P( x) is divided by ( x – r), then the remainder of this division is the same as evaluating P( r), and evaluating P( r) for some polynomial P( x) is the same as finding the remainder of P( x) divided by ( x – r).

Directrix (directrix):

@blahblah_who_cares

OpenStudy (anonymous):

Yah. Ive never learned Remainder Theorem. @Directrix

Directrix (directrix):

We'll do synthetic division.

Directrix (directrix):

P(x)=x^3+17 is divided by (x+2) P(x) = x^3 + 0x^2 + 0x + 17 divided by (x + 2)